OpenStudy (anonymous):

solve system of equations algebraically: 35x-17y=560 24x+15y=2250

OpenStudy (anonymous):

You mean using substitution? Or is it alright if we use Gaussian elimination?

OpenStudy (anonymous):

ah sorry, using substitution.

OpenStudy (anonymous):

from the graph it should be 50, 70 but I'm not getting the correct answer when solving.

hero (hero):

Write both equations in terms of y first

hero (hero):

Then set y = y, cross multiply and solve for x

OpenStudy (anonymous):

x=(570+17y)/35 (24(560+17y)/35) +15y=2250 ==> y=70 Insert y=70 in first equation (35x - 17y=560) gives you: 35x-17*70=560 ==> x= 50

OpenStudy (anonymous):

Explanation: 1. Solve for x in first equation. 2. Insert x from first equation into second equation. 3. Solver for y 4. Insert y into first equation 5. Solver for x

hero (hero):

35x-17y=560 24x+15y=2250 y = (560 - 35x)/-17 y = (2250 - 24x)/15 y = y 15(560 - 35x) = -17(2250 - 24x) 8400 - 525x = -38250 + 408x 46650 = 933x 46650/933 = x 50 = x

hero (hero):

That's a full solution